Write the equation of the line passing through (-1,4) and (4,-1). Write the equation in standard form Ax+By=C.
Answer by mananth(16949)

Write the equation of the line passing through (-1,4) and (4,-1). Write the equation in standard form Ax+By=C.
...
slope m = (-1-4)/(4+1)
m=-1
...
y=mx+b
take any point say(4,-1), m=-1
plug them
-1=-1*4+b
-1+4=b
b=3
...
equation is
y=-x+3
x+y=3
...
m.ananth@hotmail.ca